# Think Twice: Branch-and-Rethink Reasoning Reward Model

### v2

@@ Appendix A Preliminary Study on Judgment Diffusion
We follow a clear procedure to gather our results. First, for every task in our test sets, we have both the Baseline and Our Method generate their respective judgments. The next and most critical step is Token Allocation. Here, we prompt a large language model, GPT-4o-0806, to read the generated text from each model and attribute every single sentence to one of the predefined evaluation dimensions (e.g., Information Accuracy, Logical Reasoning). For our method, we combine the tokens from both the initial Adaptive Branching turn and the final Branch-Conditioned Rethinking turn. After that, we calculate the proportions by determining the percentage of total test-time compute (tokens) spent on each dimension for every instance. Finally, we average these percentage allocations across all instances within each split.
@@ B.3 Baselines
  • Scalar RMs: Conventional reward models that directly output a single scalar score without generating intermediate reasoning. We compare to SteerLM-RM-70B, Eurus-RM-7B (47), InternLM2-20B-Reward, InternLM2-7B-Reward (6), Skywork-Reward-Gemma-2-27B , Skywork-Reward-Llama-3.1-8B (19), ArmoRM-Llama3-8B-v0.1, Nemotron-4-340B-Reward (1), and INFORM-Llama3.1-70B (9).
@@ Appendix E Detailed Experiment Results
In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Tables 6, 7, and 8.

### v3

@@ Appendix A Preliminary Study on Judgment Diffusion
We follow a clear procedure to gather our results. First, for every task in our test sets, we have both the Baseline and Our Method generate their respective judgments. The next and most critical step is Token Allocation. Here, we prompt a large language model, GPT-4o-0806 (25), to read the generated text from each model and attribute every single sentence to one of the predefined evaluation dimensions (e.g., Information Accuracy, Logical Reasoning). For our method, we combine the tokens from both the initial Adaptive Branching turn and the final Branch-Conditioned Rethinking turn. After that, we calculate the proportions by determining the percentage of total test-time compute (tokens) spent on each dimension for every instance. Finally, we average these percentage allocations across all instances within each split.
@@ B.3 Baselines
  • Scalar RMs: Conventional reward models that directly output a single scalar score without generating intermediate reasoning. We compare to SteerLM-RM-70B, Eurus-RM-7B (47), InternLM2-20B-Reward, InternLM2-7B-Reward (6), Skywork-Reward-Gemma-2-27B , Skywork-Reward-Llama-3.1-8B (19), ArmoRM-Llama3-8B-v0.1, Nemotron-4-340B-Reward (1), and INF-ORM-Llama3.1-70B (9).
@@ E.1 Main Results
In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Tables 4, 5, and 6.
Table 1 presents a comprehensive performance comparison of our proposed models (BR-RMs) against the best-performing baselines across three standard benchmarks.
ScalarRMs remain competitive but plateau. ScalarRMs like INF-ORM-Llama3.1-70B achieve the strongest RewardBench score (95.1) and lead most scalar baselines in average accuracy (78.8). However, despite high scores on RewardBench, they lag behind reasoning-focused models on RM-Bench (70.9) and RMB (70.5). Smaller ScalarRMs perform relatively well, suggesting simply scaling up parameters without reasoning mechanisms brings diminishing returns.
E.2 Ablation Study
Prompting strategy BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
BR-RM (Full Model) 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
Branching Only 90.1 82.6 67.4 80.0 90.9 83.2 71.0 81.7
Unconditioned Rethink 90.6 83.9 69.3 81.3 91.3 84.4 71.9 82.5
Single-Turn Model 90.4 84.5 70.0 81.6 91.3 85.2 73.3 83.3
E.3 Reward Design Analysis
Scoring on a Scale In another experiment, we had the model output a detailed score on a scale, like from -3 to 3. This is inspired by the reward dataset, HelpSteer3, whose human-annotated preferences use a similar scale to express preference strength. During training, the reward was based on the distance between the predicted score and the ground-truth score. To process other datasets with only binary labels, we arbitrarily mapped the winning response to +2 and the losing one to -2. This approach failed, possibly due to a disconnect between the training task and the final evaluation: while training rewarded the model for predicting an exact score, our ultimate evaluation only cared about whether the model correctly identified the winning response. This means that the model had no strong incentive to learn the true meaning behind the scores; it only needed to learn if a score was positive or negative. Predicting precise scores is less beneficial when only the winner matters.
The Best Reward: Binary Reward with Format Checking. After testing these alternatives, we confirmed that the best reward design is the most straightforward. In this setup, the model produces a binary choice within a required format and receives a simple reward based on whether it is correct. Each of the other designs introduced a flaw – confusing signals, misaligned objectives, or indirect feedback – that hurt performance. This shows that a direct alignment between the task and a reward based on the correct answer provides the most effective foundation for training.
Reward BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
Our Reward 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
Removing Format Checks 90.5 84.2 71.5 82.1 91.3 85.4 73.1 83.3
Scoring on a Scale 90.0 83.5 70.1 81.2 90.8 85.1 71.7 82.5
Additional Reward for Branch 86.5 79.0 66.2 77.2 87.8 82.7 69.3 79.9
E.4 Training Data Analysis
To understand each dataset’s contribution, we conducted an ablation study, systematically removing one at a time to measure its impact on model performance. From Table 9, this analysis revealed a clear distinction between the roles of foundational and specialized datasets. For HelpSteer3, removing this foundational dataset led to a broad performance decline, with a 4.4-point drop on the RMB benchmark and a 1.5-point drop on RMBench. This highlights its critical role in establishing general understanding and factual accuracy. For Skywork, excluding our only dataset with safety data caused the most severe impact—a 5.0-point drop on the safety-focused RMB benchmark. This confirms that explicit safety training is essential for model harmlessness. For Math-Step-DPO, removing this specialized dataset predictably weakened reasoning, causing a 2.1-point drop on RMB and a 0.9-point drop on RMBench. This shows in-domain data is necessary for complex reasoning. For Code-Preference, the exclusion of this smaller dataset resulted in the most modest decrease (0.7 points on RMB and 0.5 points on RMBench). This suggests its logical reasoning contributions are partially covered by the broader datasets.
Model BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
Full training data 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
  w/o HelpSteer3 90.4 83.5 68.0 80.6 91.5 84.4 70.3 82.1
  w/o Skywork 90.6 84.2 66.1 80.3 91.3 85.1 69.7 82.0
  w/o Math-Step-DPO 91.0 84.7 70.7 82.1 91.5 85.0 72.6 83.0
  w/o Code-Preference 91.1 84.8 71.2 82.4 91.8 85.4 74.0 83.7